Please help me solve: w-c/w-d = c^2/d^2 (for w)
Answer by rothauserc(4718)

(w-c) / (w-d) = c^2 / d^2
cross multiply the fractions
d^2(w-c) = c^2(w-d)
d^2w - d^2c = c^2w - c^2d
add d^2c to both sides of =
d^2w = c^2w - c^2d + d^2c
subtract c^2w from both sides of =
d^2w - c^2w = d^2c - c^2d
w(d^2 - c^2) = (d^2c - c^2d)
w = (d^2c - c^2d) / (d^2 - c^2)
w = (cd(d - c)) / ((d - c)(d + c))
w = cd / (d + c)